### v1

# Exploring Remote Collaboration: The Impact of Avatar Representation on Dyadic Haptic Interactions in Shared Virtual Environments

Abstract
This study is the first to explore the interplay between haptic interaction and avatar representation in Shared Virtual Environments (SVEs). We focus on their combined effect on social presence and task-related scores in dyadic collaborations. In a series of experiments, participants performed the plate control task with haptic interaction under four avatar representation conditions: avatars of both participant and partner were displayed, only the participant’s avatar was displayed, only the partner’s avatar was displayed, and no avatars were displayed. The study finds that avatar representation, especially of the partner, significantly enhances the perception of social presence, which haptic interaction alone does not fully achieve. In contrast, neither the presence nor the type of avatar representation impacts the task performance or participants’ force effort of the task, suggesting that haptic interaction provides sufficient interaction cues for the execution of the task. These results underscore the significance of integrating both visual and haptic modalities to optimize remote collaboration experiences in virtual environments, ensuring effective communication and a strong sense of social presence.
Index Terms:
Avatar Representation, Haptic Interaction, Remote Collaboration
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
Virtual Environment (VE) technologies have provided unprecedented new experiences to users. In particular, Virtual Reality (VR) devices, represented by recent advances in head-mounted displays (HMDs), have become more performant and affordable, making VR increasingly accessible to the general public. In VEs, users can represent body movements through avatar embodiment[1]. Various avatar representations have been reported to alter user experiences[2]. Accurate measurement of self-body and reflection in an avatar can enhance the sense of ownership of the body[3]. Furthermore, accurate body tracking enhances the user experience in VEs[4].
Shared virtual environments (SVEs), where VE is shared with other users, allow users to share space despite each user being remotely located. SVEs are attracting attention as a new communication tool.[5]. In SVEs, visual interaction, which is crucial for users to communicate with each other[6], enhances social presence by providing the feeling of “being there”[7].
Several studies have reported that the absence of an avatar representation can reduce social presence[8, 9], a decrease in social presence can potentially lead to reduced social communication, which is a critical issue. An avatar is one of the critical elements that provide a visual modality in non-verbal communication. Moreover, remote collaboration using SVEs is expected[10]. Remote collaboration through SVEs allows you to solve tasks collaboratively without location constraints, potentially solving issues such as labor shortages.
In recent years, VR applications that integrate haptic technologies have attracted attention. Haptic feedback can provide information that cannot be conveyed visually and is an essential modality for humans[11]. The haptic modality is crucial when touching objects or moving, but it also builds a beneficial communication channel during interactions with others[12, 13]. In SVEs, haptic interaction provides various benefits to users. This means that haptic interactions influence not only physical performance but also the psychological user experience. Some studies have found that haptic interactions contribute to users feeling a higher social presence[14, 15, 16, 17, 18]. Therefore, haptic interaction is indispensable for performing physical tasks in remote collaboration.
Currently, it is still unclear whether displaying avatars is necessary for physical remote collaboration. To our knowledge, no study has investigated the impact of avatar representation on users when there is haptic interaction. Does haptic interaction alone provide sufficient social presence, or is the representation of avatars also necessary? We found no clues to this question.
Therefore, this paper compares the representation of avatars in dyadic physical remote collaboration under four conditions (both the user’s and the partner’s avatars are displayed, only the user’s avatar is displayed, only the partner’s avatar is displayed, and no avatars are displayed). Social presence, task performance, and users’ force effort were measured. The results showed that participants felt a decrease in social presence when the partner’s avatar was not displayed. Meanwhile, task performance and effort remained constant under all conditions. Therefore, participants found the social presence provided by haptic interactions to be insufficient, but the task execution was not affected by the representation of avatars.
Based on these results, we derive the necessity of avatar representation for developing physical remote collaboration. The contributions of this paper are as follows: 1) When the purpose is solely task execution in physical remote collaboration, the representation of avatars can be omitted. 2) In social communication channels, visual interaction through avatar representation is more important than haptic interaction.
I RELATED WORK
I-A Avatar representation in shared virtual environment
In social interactions, visual information plays a crucial role. Especially in SVEs provided through HMDs, users expect a bodily experience that mimics real-world movements and synchronizes with their own actions. Avatars, as a form of visual information, not only offer users a realistic representation of their bodies, but also facilitate the presence of others[19]. The presence of avatars of others, compared to audio and video communication channels, sufficiently enhances the user’s experience of emotional intimacy[6]. The representation of avatars plays a crucial role in determining user experience and task performance[20]. Gao et al. reported that the use of full-body avatars improved task performance compared to using only hands or hands and torso[21]. Moreover, Pan et al. conducted competitive and cooperative tasks in virtual environments with and without self-avatars and in face-to-face settings[10]. They found that the absence of a self-avatar during cooperative tasks led to a decrease in task scores. However, no significant differences were observed in competitive tasks. Thus, the impact of avatar representation on performance may vary depending on the context of the task.
Furthermore, avatar representation contributes to enhancing social presence. Aseeri et al. compared three conditions: no avatar, scanned avatar, and real avatar, and found that scanned avatar and real avatar demonstrated high levels of social presence[8]. In addition, complete mapping[9], full-body tracking[22], and facial expressions[23] in avatars improve social presence. In summary, representing avatars in SVE significantly affects users and contributes strongly to the formation of communication channels.
I-B Social presence and haptic interaction in shared virtual environment
Haptic interaction allows users to share forces among themselves, enabling them to perform physical remote collaboration. Physically connected dyads form communication channels through the exchange of forces, allowing them to receive information from their partner [12]. Haptic interactions offer various physical benefits to dyads: improvement in time to motor learning [24, 25], enhancement of performance [26], maintenance of balance [27], and negotiation of strategies [12].
Haptic interactions also provide user experience benefits beyond physical advantages. Several studies have investigated the impact of haptic interactions in SVEs by toggling the presence and absence of such interactions among dyads [14, 15, 16, 17, 18]. Chellali et al. conducted a task in which dyads trained in injection procedures using a Follow-Follower system and demonstrated that haptic interactions improved performance and co-presence[15]. Additionally, Sallnas et al. found that exchanging objects through haptic interactions enhanced social presence[14]. In summary, haptic interactions provide both physical and psychological benefits to users.
Both visual interaction through avatar representation and haptic interaction have been confirmed to enhance social presence and task performance. The question arises: Can social presence and task performance be maintained through only haptic interactions when avatar representation is lost? To date, no studies have focused on both avatar representation and haptic interactions simultaneously. Our study is the first to address this question.
II EXPERIMENTAL SETUP
We developed an experimental environment for a physical remote collaboration task as shown in Fig. 1. The SVE displayed a plate and avatars representing the participants’ body movements. The dyes manipulated the plate through a haptic device, performing cooperative tasks.
Figure 1: Experimental setup. (a) The shared virtual environment. (b) Participants could perform the cooperative task through the independent haptic interface. (c) Participant install the HMD and see the SVE. (d) The plate control task.
II-A Shared virtual environments
Our developed SVE was created using Unity 2022.3.23f1. In this SVE, a plate was fixed in a floating position at a height of 1 m, surrounded by avatars. Dyad’s avatars were positioned face-to-face across the plate. The plate is a square with each side 0.4 m. Participants accessed the SVE using the Meta Quest 3, which tracked the position of participants’ heads, wrists, and each joint of the fingers, applying these positions in real time to the avatars’ actions. Data from Meta Quest 3 were captured using the Meta XR Core SDK v63.0.0. The movements of the avatar’s body parts not derived from captured positions were determined using Final IK. Inverse Kinematics (IK) techniques can cause errors, which can cause a feeling of discomfort for participants[28]. However, since the participants moved within a limited range in this study, these errors were not significant enough to affect the experiment and were not be considered.
The avatars used were generic with no gender or distinctive features, reflecting only the participants’ height as physical information. This choice was made to eliminate any impact on the facial expressions or gender of the participants. In fact, the force exerted on an avatar has been reported to decrease when it appears to be female[29]. Furthermore, participants interacting with expressive avatars feel stronger social presence and attractiveness and show superior task performance compared to those interacting with non-expressive avatars[30]. We needed to remove these variables to study primitive effects.
II-B Haptic Interface
Haptic interactions between dyads were bilateral. One of these haptic interfaces is provided for each participant and operates independently. The haptic interface featured two motors and had a gimbal mechanism that could rotate along the roll and pitch axes. Thus, participants were capable of two degrees of freedom of movement. The motors used were Mitsubishi Electric’s HG-MR43, and to maintain high backdrivability, they were driven directly. Direct drive is often adopted in the research of physical Human–Human Interaction (pHHI) [31, 13]. Extending from the gimbal mechanism was a rod that could be gripped with both hands. The length of the rod from the left to the right was 0.4 m, and the length to the gimbal mechanism was 0.2 m, perfectly corresponding to the length of the plate. The gimbal mechanism was fixed at a height of 1 m and the participants operated it through the haptic interface in a standing position.
II-C Haptic interaction between dyad
To provide haptic interaction to the dyad, the study applied acceleration-based admittance control to the haptic interfaces, allowing bilateral driving [32]. This form of control, suitable for building haptic interactions, allows precise transmission of forces among participants [33]. A disturbance observer was used to implement acceleration-based control [34], which suppresses external forces in the control system, aligning the command acceleration with the controlled acceleration. Furthermore, a reaction torque observer was employed to measure participants’ forces without a force sensor [35].
Admittance control defines how a system responds to human operator forces through its dynamics, allowing designers to arbitrarily design these dynamics. This capability allows haptic interactions to be provided as if the dyads are jointly manipulating a single object in reality, a method employed in pHHI studies [12]. In this research, the dynamics of the plate were shared among the dyads.
Here, the dyad’s force inputs through the haptic interfaces produced the forces that actuated the plate:
${\bm{\tau}^{plt}}=\left[\begin{matrix}\tau_{1}^{p}+\tau_{2}^{p}\\ \tau_{1}^{r}+\tau_{2}^{r}\end{matrix}\right],$ (1)
where $\tau_{1}$ and $\tau_{2}$ represent each participant’s force input, with p and r indicating the pitch and roll axes, respectively.
The dynamics of the plate are expressed as:
${\bm{\tau}^{plt}}=M\ddot{\bm{\theta}}^{plt}+D\dot{\bm{\theta}}^{plt},$ (2)
where $\ddot{\bm{\theta}}^{plt}$ and $\dot{\bm{\theta}}^{plt}$ represent the plate’s angular acceleration and velocity vectors, respectively, integrated to obtain the plate’s angular vector $\bm{\theta}^{plt}$. $\ddot{\bm{\theta}}^{plt}$, $\dot{\bm{\theta}}^{plt}$ and $\bm{\theta}^{plt}$ are
$\ddot{\bm{\theta}}^{plt}$ $=\left[\begin{matrix}\ddot{\theta}^{plt}&\ddot{\phi}^{plt}\end{matrix}\right]^{T}$
$\dot{\bm{\theta}}^{plt}$ $=\left[\begin{matrix}\ddot{\theta}^{plt}&\ddot{\phi}^{plt}\end{matrix}\right]^{T}$ (3)
${\bm{\theta}}^{plt}$ $=\left[\begin{matrix}\theta^{plt}&\phi^{plt}\end{matrix}\right]^{T}.$
$\theta$ and $\phi$ stand for pitch and roll axes angle, respectively. These dynamics, with $M$ and $D$ representing the plate’s mass and viscous coefficient, respectively, used values of 3.28 and 1.5 in the experiment.
II-D Plate control task
A ball and a target were placed on the plate. The dyads were instructed to position the ball on the target. This task, which we call the plate control task, required precise control and coordination.
The trajectory of the target followed a rose curve, which moved according to the progression of time. The movement of the target was calculated using Equation (4).
${\bm{x}^{tgt}}=r\left[\begin{matrix}\sin\left(\frac{4\pi\cdot t\operatorname{fmod}(t^{per})}{dt}\right)\cos\left(\frac{2\pi\cdot t\operatorname{fmod}(t^{per})}{dt}\right)\\ \sin\left(\frac{4\pi\cdot t\operatorname{fmod}(t^{per})}{dt}\right)\sin\left(\frac{2\pi\cdot t\operatorname{fmod}(t^{per})}{dt}\right)\end{matrix}\right].$ (4)
Here, ${\bm{x}^{tgt}}=\left[\begin{matrix}x^{tgt}&y^{tgt}\end{matrix}\right]^{T}$ represents the position vector of the target on the plate, indicating its x and y axes position. $t$is the time, and $t^{per}$ is the period of the rose curve from start to finish, designed to be 20 seconds in this study. $r$ is the radius of the trajectory, set to 0.2 m. Dyads were not shown the trajectory visually during the experiment but memorized it during practice sessions.
The ball moved according to the plate’s inclination, governed by the dynamics applied in Equation (5).
${\bm{x}^{bll}}=mg\sin{\bm{\theta}^{plt}}-\eta\dot{\bm{x}}^{bll}.$ (5)
Here, ${\bm{x}^{bll}}=\left[\begin{matrix}x^{bll}&y^{bll}\end{matrix}\right]^{T}$ is the position vector of the ball on the plate. $\dot{\bm{x}}^{bll}$ is the velocity vector of the ball. $m$ is the ball mass, $g$ is the gravitational acceleration, $\eta$ is the viscosity coefficient, with values set to 0.014, 9.8, and 1.5, respectively, in this study. $m$ and $\eta$ are critical parameters determining the movement of the ball relative to the tilt of the plate, significantly influencing the difficulty of the task. These parameters were chosen experimentally to set the difficulty of the task.
II-E Co-localization
Synchronizing the positions of the haptic interface and the VE has significant implications on task performance and user perception, as noted in previous studies [36, 37]. Ensuring that the haptic interface and the plate displayed within the SVE were precisely aligned was crucial. To facilitate this, a Meta Quest 3 controller was attached to the top of the haptic interface, serving as a landmark to provide exact positional data.
Before starting the experiments, participants were calibrated to align the landmark with the plate’s position in the SVE. This setup allowed participants to manipulate the SVE plate through the haptic interface without any sense of discrepancy or discomfort.
The landmark was also used to share the positions of avatars. When a participant grasped the plate, it was essential that the avatar of their partner appeared to be grasping the plate as well, enhancing the sense of collaboration and presence within the SVE. The avatar’s position was captured from the landmark’s local coordinate system and transmitted to other devices, ensuring that the dyad could share the space around the plate completely, irrespective of the interface’s physical location. The avatar position data were updated and transmitted every 10 milliseconds, allowing participants to focus on their tasks without experiencing any perceivable delay. This technology facilitated a seamless integration of physical and virtual interactions, crucial for maintaining high levels of engagement and performance in collaborative virtual tasks.
III EXPERIMENT
This paper investigates the impact of avatar representation in remote physical collaboration. We hypothesized that haptic interaction alone could sustain social presence among dyads, even without the visual representation of avatars. To test this, we conducted an experiment in which dyads performed a plate control task through haptic interaction while varying the avatar representation of the participants and their partners. Previous research indicates that the absence of an avatar significantly reduces social presence [8]. Consequently, the variation in avatar representation was limited to visible or invisible in our experiment.
III-A Experimental design
Figure 2: Participant’s view on each condition of avatar representation. From left to right, are ALL, ALONE, OTHER, and NONE.
To assess the impact of avatar representation, we compared four conditions of avatar representation:
  • ALL: Both participant’s and partner’s avatars were displayed in the virtual environment. It assumed that both users had body tracking devices.
  • ALONE: Only participant’s avatar was displayed. Although both users had body tracking devices, avatar-related data was not transmitted to the partner. This condition could save communication bandwidth, ensuring reliable transmission of haptic information.
  • OTHER: Only the partner’s avatar was displayed.
  • NONE: No avatars were displayed. This simplifies and reduces the cost of SVE development as no body-tracking device is required.
The conditions ALL, ALONE, and NONE are those anticipated in real environments. On the other hand, OTHER has been newly added based on our previous research [38], but it is a condition that does not occur in reality (as only the partner receives body tracking information, and no avatar is displayed on one’s own screen). However, we were interested in the effects that occur when self-avatar is not displayed. If self-avatar alone is not displayed in the SVE, can participants maintain their sense of social presence? In addition, each condition was applied simultaneously to both participants.
III-B Research hypothesis
Haptic interaction and visual interaction through avatar representation contribute to providing the presence of a partner. Even without avatar representation, participants may still feel the presence of the other through haptic interaction, and the sense of social presence may not be lost. On the other hand, we speculate that dyads will work hard to achieve the objectives of the task, regardless of the conditions. Therefore, we propose the following hypotheses:
  • H1: Haptic interaction continues to provide the presence of a partner, even when avatars are not displayed. Thus, the social presence remains the same across all conditions.
  • H2: Regardless of avatar representation, the dyads will continue to achieve the best performance possible. Thus, the performance is consistent across all conditions.
  • H3: Changes in avatar representation can psychologically affect participants and alter the intensity of their efforts. Thus, there is a variation in the intensity of the effort between conditions.
III-C Participants
28 participants were included in this experiment, all university students. Participants provided informed consent according to the Declaration of Helsinki before the experiment. The experiment was approved by the Human Life Ethics Committee of Tokyo Denki University. Among the participants, 24 were right-handed and four left-handed. The group included 24 males and four females aged 21 to 28 ($M=22.96$, $SD=2.08$).
Participants were randomly paired into 14 dyads, with relationships ranging from friends and acquaintances to strangers. The visibility of a friend’s face during a collaborative task has been reported to influence performance [39]. However, the absence of visibility of the face does not affect the results, even with close friends [40]. Therefore, the relationship between partners was not considered in this experiment. All participants were first briefed on the experiment’s objectives, followed by an explanation of the equipment used.
III-D Procedure
Initially, participants provided personal information (biological sex, age, height, and dominant hand) to collect physical data. The participants then practiced alone to familiarize themselves with the plate control task, SVE, and HMD. They wore the Meta Quest 3 and calibrated it to synchronize the haptic interface with the virtual environment, verifying that their avatar tracked their body, hands, and fingers accurately. During practice, participants could ask the experimenter questions about the plate control task without restrictions. The practice continued until participants declared comfortable with the plate control task, with no time limits set by the experimenter.
After practice, each dyad collaboratively performed the plate control task for 80 seconds. The avatar representation condition was applied randomly from four conditions, and visual interaction between the dyads was presented. During the experiment, verbal communication between the dyads was prohibited by the experimenter. Following the plate control task, participants removed the Meta Quest 3 and completed a questionnaire as instructed by the experimenter (details in III-E1). The questionnaire was displayed on a computer. After completing the questionnaire, the participants were given at least a two-hour break to consider fatigue and memory of their responses. After the break, participants performed the plate control task under a different condition and completed the questionnaire again. The experimenter conducted the plate control task and questionnaire four times with each pair to compare the four conditions.
III-E Measures
A systematic review on the effects of avatars and agents during the use of HMD highlights that avatars can be effective or ineffective, depending on the task and situation [19]. They also point out that combining subjective evaluations with numerical data enhances the validity of research findings. Bailenson et al. argue that subjective measures, self-reported, depend on the participant’s ability to accurately assess their experiences [41]. Subjective measures are susceptible to participant bias, which can affect reliability. However, objective measures can detect responses to digital representations, such as personal distance, which are not captured by self-reports, suggesting higher sensitivity. Thus, in this paper, user experience (social presence) is subjectively measured, while task-related metrics (performance and effort) are objectively measured.
III-E1 Subjective measures
Social presence is the psychological sense that another person is “being there” when using media or communication technologies [42]. This concept delivers the intimacy and immediacy of real interactions despite physical separation. A higher richness of information in communication media can create a stronger sense of social presence. To deepen understanding of how people feel connected in virtual environments, social presence is often used [7].
To measure social presence, this study used the Networked Minds Social Presence Questionnaire (NMSPQ) [43, 44]. The NMSPQ includes five sub-scales: Co-presence, Perceived Attentional Engagement, Perceived Emotional Contagion, Perceived Comprehension, and Perceived Behavioral Interdependence. Since the interaction in our task was limited to non-verbal communication, participants had no opportunity for conversation and could not comprehend their partner’s thoughts. In addition, the general avatars used were expressionless and lacked personality, which prevented the transmission of emotions to the partner. For these reasons, Perceived Comprehension and Perceived Emotional Contagion were removed from the questionnaire.
Consequently, only the relevant sub-scales of Co-presence, Perceived Attentional Engagement, and Perceived Behavioral Interdependence were retained. The questions were adjusted to align with the experimental context, and the questionnaire comprised 20 items. Responses were collected on a 7-point Likert scale ranging from “Strongly disagree” to “Strongly agree.”
III-E2 Objective measures
In this paper, performance and effort were used as objective measures. The objective of the plate control task was to move a ball on a target, which required the dyads to collaborate to decrease the distance between the target and the ball. We decided to measure performance by calculating the root-mean-squared error (RMSE) between the target and the ball. Performance is defined as follows:
$P=\frac{\sum_{k=1}^{N}\left|\left|{\bm{x}_{k}^{tgt}}-{\bm{x}_{k}^{bll}}\right|\right|}{\sqrt{N}}.$ (6)
Here, $k$ represents the number of steps ($k=1\ldots N$), corresponding to the number of data points captured from the haptic interface. Note that a lower value of $P$ indicates better performance.
Effort was also measured. Although participants applied force to move the plate, not all applied force effectively contributed to the movement. According to Newton’s laws, when two equal forces are applied in opposite directions to an object, they cancel out, and the object does not move. In contrast, if forces are applied in the same direction, the object moves. Thus, the difference between the total forces applied to the object and the forces that actually moved the object represents the effort not contributing to the movement, considered as effort here. Although effort does not directly impact performance, it is considered crucial for haptic communication used by dyads to negotiate for achieving high performance.
The total force applied to the plate is given by:
${\bm{\tau}^{tot}}=\left[\begin{matrix}\left|\tau_{1}^{p}\right|+\left|\tau_{2}^{p}\right|\\ \left|\tau_{1}^{r}\right|+\left|\tau_{2}^{r}\right|\end{matrix}\right].$ (7)
From which the definition of Effort follows:
$F^{eff}=\frac{1}{N\cdot L}\sum_{k=1}^{N}\left[\begin{matrix}1&1\end{matrix}\right]\left({\bm{\tau}_{k}^{tot}}-\left|{\bm{\tau}_{k}^{plt}}\right|\right)$ (8)
Here, $k$ represents the number of steps ($k=1\ldots N$), $L$ is the half length of the plate used to convert torque into force. This measure helps to understand the extent of unproductive effort exerted by participants during the task.
IV RESULTS
IV-A Subjective measures
[…]
TABLE I: Post Hoc Comparisons - Social Presence
SP CP PAE PBI
MD pbonf MD pbonf MD pbonf MD pbonf
ALL ALONE $1.148$ $<$ .001 $1.612$ $<$ .001 $1.179$ $<$ .001 $0.500$ $0.147$
OTHER $0.489$ $0.062$ $0.728$ $0.035$ $0.304$ $1.000$ $0.357$ $0.633$
NONE $1.448$ $<$ .001 $2.143$ $<$ .001 $1.071$ $<$ .001 $0.899$ $<$ .001
ALONE OTHER $-0.659$ $0.004$ $-0.884$ $0.005$ $-0.875$ $0.001$ $-0.143$ $1.000$
NONE $0.300$ $0.669$ $0.531$ $0.250$ $-0.107$ $1.000$ $0.399$ $0.427$
OTHER NONE $0.959$ $<$ .001 $1.415$ $<$ .001 $0.768$ $0.005$ $0.542$ $0.091$
Note. P-value adjusted for comparing a family of 6. MD stands for Mean Difference.
SP, CP, PAE and PBI stand for Social Presence, Co-presence, Perceived Attentional Engagement
and Perceived Behavioral Interdependence, respectively.
Figure 3: Subjective measures of Social Presence (NMSPQ). From left to right, Social Presence, Co-presence, Perceived Attentional Engagement and Perceived Behavioral Interdependence.
IV-B Objective measures
Figure 4: Comparison of task performance of the plate control task.
Figure 5: Comparison of effort of the plate control task.
[…]
Given this nonnormality, the Friedman test, a nonparametric equivalent to repeated measures ANOVA, was performed to check for differences across the conditions ALL, ALONE, OTHER, and NONE in terms of Performance and Effort. The results showed that there were no significant differences in either performance or effort under the conditions. Comparisons of performance and effort under the different conditions are illustrated in Fig. 4 and Fig. 5, respectively.
Figure 6: Comparison of the temporal changes in normalized performance.
Additionally, changes in performance were observed over time. Fig. 6 shows a comparison of the temporal changes in performance. Here, a ’Quarter Period’ represents the average of each quarter of a rose curve, with the period of the rose curve being 20 s (i.e. the one-quarter period is 5 s). The plate control task was conducted for 80 seconds, so the graph displays 16 quarter periods. Performance for each quarter period was normalized based on performance in the first quarter period.
As a result, it was confirmed that the dyads improved their performance compared to the first quarter period under ALL, ALONE, OTHER and NONE conditions (lower values indicate better performance). However, no clear differences were detected between the conditions ALL, ALONE, OTHER and NONE. This suggests that while performance improved during the cooperative task, the specific condition did not significantly influence the rate or extent of improvement.
V DISCUSSION
V-A Subjective measures
The survey revealed significant differences in all measures of Social Presence, Co-presence, Perceived Attentional Engagement, and Perceived Behavioral Interdependence. Notably, significant differences were found in comparisons of the co-presence sub-scale, except between the ALONE and NONE conditions. These results led to the rejection of the hypothesis ${\bf H1}$.
The ALL condition yielded the highest scores, indicating a high level of social presence, which was not unexpected. It is well documented that the presence of a partner’s avatar increases social presence, as demonstrated through various studies [8, 22, 9, 19]. Our findings reinforce the importance of displaying a partner’s avatar in the SVE.
Social presence scores in the OTHER condition were the next highest after ALL, but ALONE and NONE showed lower scores compared to ALL and OTHER. This suggests that having a self-avatar is less critical for maintaining a high social presence than seeing the partner’s avatar. Participants were constantly engaged in collaborative tasks that involved plate handling during the experiments, a constant factor in all avatar conditions. It is clear that the partner’s presence was sensed through haptic interaction regardless of the avatar’s visibility. However, the absence of the partner’s avatar reduced the social presence. Haptic feedback plays a vital role in SVEs, and several studies have reported that it improves social presence [14, 15, 16, 17, 18]. These results imply that visual representation of a partner, such as avatars, might be more critical than haptic interaction for enhancing social presence.
Next, we focus on the perceived behavioral Interdependence. Our results indicate that the ALL condition scored significantly higher than NONE. Perceived Behavioral Interdependence measures the extent to which users’ actions affect and are affected by their partners. Thus, participants felt their influence on or influence from the partner was more significant when avatars were visible. This finding is supported by studies such as those by Heidicker [9] and Yoon [45]. However, in our study, the dyads were constantly provided with haptic feedback, and their influence on each other always existed between the dyads, regardless of avatar visibility. Yet, despite this, avatars influenced participants, potentially feeling monitored by their partners (similar to how one might behave more cautiously when a camera is on during a web conference). This observation suggests that it may be possible to choose to omit the display of avatars in order to reduce the mental burden on users.
V-B Objective measures
The measurements of Performance and Effort revealed no significant differences based on avatar representation. Despite visual limitations, the dyads continued to demonstrate the level of performance that they could achieve in the plate control task. Thus, hypothesis ${\bf H2}$ was supported. Various studies have investigated the impact of avatar representation on task performance, yielding mixed results [21, 10, 45, 30]. Yoon et al. explained that the importance of avatars varies between user-centered situations (where the main focus is on collaboration among users) and task-centered situations (where the main focus is on completing the task through collaboration) [45]. Our plate control task, which could be completed irrespective of avatar representation, falls into the category of task-centered situations, and the consistent performance across different avatar conditions supports their claim.
Moreover, two studies that examined the impact of visual interaction in remote collaboration tasks found that showing partner manipulation information as cursors on the screen improved the performance of dyad work [46, 47]. The tasks in their studies required estimating the partner’s maneuvers, where visual information about the partner was crucial for task completion. Our plate control task involved coordinating the manipulation of a rigid plate, and the information provided through haptic interaction may have been sufficient to estimate the partner’s actions. If the task had involved moving a table, where hand positioning is critical, avatar representation might significantly impact performance, especially when the partner’s physical information plays a crucial role in task completion.
In terms of effort, hypothesis ${\bf H3}$ was rejected. Dyads did not alter their effort despite changes in visual information. For physically connected dyads, the forces between partners play a crucial role, especially in negotiating and integrating intentions [12, 13]. Groten et al. demonstrated that haptic information provides a valuable channel for integrating intentions between dyads in co-performed haptic tasks. Roche et al. supported this, explaining that when dyad intentions differ, the forces exerted increase, and haptic information is utilized in negotiation. These findings suggest that the presence of an avatar did not influence users’ intentions in performing tasks. This could be because the avatars used in our study were generic and lacked gender, expressions, gaze direction, and physical cues that could significantly influence interactions. Indeed, studies have shown that avatar appearance, such as gender, can influence how much force is perceived as appropriate; for instance, interactions with female avatars might result in gentler forces compared to male avatars [29]. Therefore, using avatars with a higher level of detail could potentially affect haptic interactions.
V-C Generalization
We develop an argument that integrates subjective and objective measures. Although changes in social presence were observed across the four avatar conditions, Performance and Effort remained unchanged.
The first key finding is that for dyads physically connected by haptic interaction, avatar representation does not impact performance or the quality of haptic interactions. This insight could significantly influence the development of physical remote collaboration applications. Implementing tracking systems for such systems may no longer be essential. Typically, developing virtual environments that display avatars requires cameras or sensor-based tracking systems, which are subject to constraints such as location and lighting [3]. Additionally, communication and rendering technologies are needed, which contribute to the cost of system construction in both hardware and software [19]. Thus, our findings could lead to reductions in various costs associated with these systems.
The second key finding is that the enhancement of social presence provided by avatar representation cannot be substituted by haptic interaction alone. This finding is particularly relevant for the entertainment sector, where user experience is crucial. Presenting avatar representations in real time to users in a SVE is sufficient to enhance the experience of emotional closeness, much like audio and video [6]. This suggests that avatars play a starring role in social communication channels. Our findings indicate that, on the contrary, haptic interactions play a supporting role. Consequently, avatar representation might be more important than haptic interaction for enhancing user experience, emphasizing the significance of visual elements in the perception of social presence and interaction quality.
This distinction highlights the critical role avatars play in the user’s emotional and psychological engagement within virtual environments, suggesting a strategic focus on visual representation to enhance social presence effectively, particularly in applications where the social dimension is predominant.
VI LIMITATION AND FUTURE WORK
VI-A Level of Detail on avatar representation
The avatars used in this study were generic and lacked detailed physical information, which was intentional for a preliminary investigation. However, the level of detail (LoD) in avatar representation could significantly impact user experience.
The Proteus effect explains how changes in self-avatar appearance can influence user behavior and cognition [48]. Use of muscular avatars has been reported to lead to reduced perceived effort and altered actions [49, 50]. Furthermore, the use of personalized avatars can improve body ownership and immersion in virtual spaces [3, 51].
In SVEs, avatar expression also has a significant impact on users. A study by Kang et al. reported that a low LoD in facial expressions of avatars can diminish social presence [23]. Conversely, interacting with expressive avatars can enhance social presence [30, 52, 53].
This suggests that varying the appearance and expressions of avatars could extend the discussions presented in this paper. Typically, cameras are used to apply facial expressions to avatars, but integrating them with an HMD presents challenges [54]. To address this, facial tracking devices that can be mounted on HMDs have been developed, allowing expressions to be represented even while wearing an HMD [55, 56]. Recently, Apple has developed a personalized avatar system, known as Persona, which can express facial details using their Vision Pro device [57].
In future studies, we plan to use these advanced devices to expand our investigation and further explore how enhanced avatar realism and expressiveness impact user experience in SVEs. This approach may reveal deeper insights into the behavioral and psychological effects of high-fidelity avatars, potentially guiding the design of more engaging and effective virtual reality experiences.
VI-B The impact on group size
In this paper, participants were limited to dyads, but there is potential to expand the research to groups larger than three people. Generally, as the group size increases, the dynamics of the group become more pronounced, and various psychological effects occur. For instance, larger groups might see an increase in free riders and a decrease in individual proactivity and performance. This decrease can be explained by phenomena such as the Ringelmann effect or the bystander effect [58, 59].
In our investigation, we found only one study that investigated the impact of changes in group size on social presence within an SVE [60]. Their research involved tasks where 2–4 person groups used HMDs to communicate, examining social presence. They found that smaller groups tend to engage more closely with each other, experiencing a higher social presence. In contrast, larger groups might feel a certain distance in their interactions with others, which can affect the sense of social presence.
Given these findings, it is clear that increasing group size could have significant impacts, but actual effects would need to be investigated. Fortunately, our developed system supports collaborative work for up to four people, allowing future studies to explore the effects of group size. This exploration could significantly contribute to understanding how social presence is mediated by group dynamics in virtual environments, potentially influencing the design of virtual reality systems and applications for larger groups.
VI-C Collaboration task
The plate control task developed and used in this paper was a task that involved coordinated manipulation of a rigid body. This task required technical and precise adjustment of the force. Future research using different types of haptic sharing tasks may develop the discussion further.
Maroger et al. conducted research on the task of carrying a table in dyads. Carrying a table is a routine and seemingly simple context, but they showed that the strategies for carrying a table vary by dyad and are complex [61]. Sawers et al. conducted research with an interesting task of dancing in dyads [62]. Their study demonstrated the potential for small interaction forces to convey movement goals.
Thus, the impact on user experience may vary depending on the type of haptic sharing task. Future research is expected to expand the discussions of this paper through various haptic sharing tasks.
VII CONCLUSION
The purpose of this paper was to investigate how avatar representation affects social presence and collaborative tasks in a Shared Virtual Environment (SVE) with haptic interaction. Four conditions of avatar representation were prepared: both the participant’s and the partner’s avatars are displayed (ALL), only the participant’s avatar is displayed (ALONE), only the partner’s avatar is displayed (OTHER), and no avatars are displayed (NONE). The experiment involved a plate control task with haptic interaction, which provided consistency among conditions. The analysis was carried out through subjective measures of social presence and objective measures of performance and effort. Social presence scored higher in ALL and OTHER conditions and lower in the ALONE and NONE conditions. Therefore, although haptic interaction provided the presence of a partner, the presence of the partner’s avatar was suggested to be important for enhancing social presence. Performance and effort remained consistent regardless of avatar representation. This suggests that avatar representation does not affect haptic interaction. From these results, we present the following contributions: 1) Avatar representation does not impact performance or effort of haptic interaction, therefore avatars can be omitted when the sole purpose is physical collaboration. 2) Visual interaction through avatars plays a leading role in social communication channels, but haptic interaction can only play a supporting role.
We hope that the findings of this paper will aid in the development of physical remote collaboration through SVEs.
Acknowledgments
This work was partially supported by; Research Institute for Science and Technology of Tokyo Denki University Grant Number Q23D-07 and Q24D-01 / Japan and JSPS KAKENHI Grant Number 22H01455.